OpenStudy (anonymous):

A triangle has coordinates A (1, 5), B (-2, 1) and C (0, -4). What are the new coordinates if the triangle is dilated with a scale factor of 1/5?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

I might be wrong but I think you should multiply each number by 1/5 to get the new coordinates

OpenStudy (anonymous):

Each number?

jigglypuff314 (jigglypuff314):

yes so A (1, 5) --> A' ( (1*1/5), (5*1/5))

OpenStudy (anonymous):

I'm still lost. What would that be?

jigglypuff314 (jigglypuff314):

1* 1/5 = 1/5 5 * 1/5 = 1 so new A (1/5, 1)

OpenStudy (anonymous):

Hmm.. I see where your going but I don't feel like that's it.

jigglypuff314 (jigglypuff314):

hmm? why not? :/

OpenStudy (anonymous):

Actually, I'll take your word for it. You know maths WAY better than I do!

jigglypuff314 (jigglypuff314):

that might not be necessarily true, but okay :P you would then do the same with the other points

OpenStudy (anonymous):

Got it. I'll go ahead and submit and see what my teacher says. Thank you so much! c:

jigglypuff314 (jigglypuff314):

good luck ^_^